# Empowering Embodied Manipulation: A Bimanual-Mobile Robot Manipulation Dataset for Household Tasks

### v1

@@ Abstract
As Embodied AI advances, it increasingly enables robots to handle the complexity of household manipulation tasks more effectively. However, the application of robots in these settings remains limited due to the scarcity of bimanual-mobile robot manipulation datasets. Existing datasets either focus solely on simple grasping tasks using single-arm robots without mobility, or collect sensor data limited to a narrow scope of sensory inputs. As a result, these datasets often fail to encapsulate the intricate and dynamic nature of real-world tasks that bimanual-mobile robots are expected to perform. To address these limitations, we introduce BRMData, a Bimanual-mobile Robot Manipulation Dataset designed specifically for household applications. The dataset includes 10 diverse household tasks, ranging from simple single-arm manipulation to more complex dual-arm and mobile manipulations. It is collected using multi-view and depth-sensing data acquisition strategies. Human-robot interactions and multi-object manipulations are integrated into the task designs to closely simulate real-world household applications. Moreover, we present a Manipulation Efficiency Score (MES) metric to evaluate both the precision and efficiency of robot manipulation methods. BRMData aims to drive the development of versatile robot manipulation technologies, specifically focusing on advancing imitation learning methods from human demonstrations. The dataset is now open-sourced and available at https://embodiedrobot.github.io/, enhancing research and development efforts in the field of Embodied Manipulation.
@@ 1 Introduction
Recent advancements in large language models (LLMs) [1, 2, 3, 4] and multimodal large language models (MLLMs) [5, 6, 7] have significantly advanced the field of Embodied Manipulation, enhancing the capacity for robots to interact with 3D physical world in a more human-like manner. This technology aims to endow robots with the ability to perform manipulation tasks within household environments, mirroring human dexterity and adaptability. Household tasks present a substantial challenge due to their diversity and the complex interactions they entail with various objects. Consequently, the development of bimanual-mobile robots, which combine the manipulation capabilities of humanoid robots with autonomous mobility, is crucial. These robots must learn and adapt to a wide array of tasks, demanding sophisticated methods for versatile robot manipulation.
Comprehensive robot manipulation datasets [8, 9, 10, 11] are critical for advancing Embodied Manipulation. These datasets are essential for developing and refining algorithms that enable robots to perform complex tasks with human-like dexterity. However, existing datasets primarily focus on single-arm robotic tasks, such as RT-1 [12], RH20T [13], and RT-X [8], which are limited to basic object grasping and placement. This narrow focus fails to address the complexities and coordination demands of bimanual-mobile manipulations. Although some datasets include dual-arm manipulations, they often lack mobility features [14] or are limited to single-sensor setups [11]. This restriction hampers their ability to comprehensively represent the complex tasks that bimanual-mobile robots must perform in diverse household environments.
@@
To overcome these limitations, we introduce BRMData, a novel dataset meticulously designed to advance the development and evaluation of bimanual-mobile robots for household tasks. BRMData captures the intricacies of real-world tasks through human demonstrations, offering a true reflection of the dexterity and interaction complexity required. As depicted in Figure 1, this dataset comprises 10 diverse manipulation tasks, including single-object pick-and-place manipulation, dual-arm coordination, human-robot interactions, and mobile manipulation. These tasks are specifically selected to incrementally challenge and develop robot manipulation capabilities. Furthermore, we propose a novel Manipulation Efficiency Score (MES), which considers both the success rate and the efficiency of robot manipulation methods. This metric provides a more comprehensive evaluation, ensuring that developed manipulation algorithms not only complete tasks effectively but also in a timely manner. In summary, we present the BRMData dataset with the following key features:
Task Diversity: BRMData encompasses a broad spectrum of manipulation tasks, including single-arm and dual-arm tasks, as well as tabletop and mobile manipulations. This diversity enhances the adaptability of robot learning algorithms across various real-world scenarios.
Progressive Difficulty: The dataset features tasks with escalating complexity, from single-arm object grasping to dual-arm coordination, human-robot interactions, and multi-object pick-and-place tasks. This progression facilitates diverse robot learning methods.
Comprehensive Sensory Data: Employing RGBD camera sensors with varied frame rates, multi-view, and depth-sensing data, BRMData provides essential sensory information for testing the robustness of robot learning methods.
@@ 2 Related Work
2.1 Robot Learning
In the realm of robot learning, the utilization of human demonstrations [15, 16, 17, 18, 12] as a means to imbue robots with manipulation capabilities has garnered significant attention. Robomimic [19] conducts a comprehensive examination of six offline learning algorithms across various simulated and real-world manipulation tasks, delves into the nuances of learning from offline human demonstrations for robot manipulation. DayDreamer [20] enables robots to learn from minimal interaction by planning within a learned world model, which significantly reduces the trial-and-error process typically required in physical settings. CALVIN [21] offers environments for policy learning under language instructions for long-horizon robotic manipulation tasks, emphasizing the importance of comprehensive datasets in enhancing robot-environment interactions. Similarly, DROID [10] provides extensive data for robot manipulation, underscoring the necessity of testing and optimizing algorithms for dynamic, real-world applications. Moreover, ARMBench [22] and PLEX [23] have expanded the application scope by incorporating not only basic manipulation tasks but also more complex scenarios and diverse operational strategies.
@@ 2.2 Robot Manipulation Datasets
3 BRMData Dataset
According to the comparison provided in Table 1, BRMData stands out due to its comprehensive coverage of bimanual-mobile manipulations. Through the integration of dual-arm coordination and cross-sensor data collection, BRMData offers an enriched training environment unparalleled by existing representative datasets. This section describes the methods employed in the data collection and details the composition of manipulation tasks.
Table 1: Comparison to existing datasets for robot manipulation.
Dataset Human Demonstration Movable Dual-arm Depth-sensing Multi-view Torque Cross-sensor
MIME [15] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$
RT-1 [12] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$
RH20T [13] $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$
BridgeData V2 [25] $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$
ALOHA [14] $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Mobile ALOHA [11] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
BRMData (Ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
3.1 System Setup
Figure 2: Illustration of hardware details.
3.2 Data Collection and Processing
@@
The data collection strategy for the BRMData dataset is meticulously designed to comprehensively capture a wide array of manipulation tasks. As shown in Figure 3, both the left and right wrist-mounted RGBD cameras and the middle camera record RGB and depth images at a resolution of $640\times 480$. The integration of these three views provides a multi-perspective visual input, essential for constructing a detailed 3D representation of the operational environment. For specific tasks such as the Bottle Pick and Bottle Handoff, image data are captured using Orbbec Dabai cameras operating at a frequency of 30Hz. For more dynamic manipulation tasks, Intel Realsense D435 cameras, which captures data at a higher frequency of 60Hz, is employed to accommodate the quick movements and frequent data updates required. Each manipulation task within the dataset includes the collection of 50 action trajectories, gathered at intervals that are specifically tailored to the complexity and specific requirements of each task scenario. To enrich the diversity of the dataset, each manipulation task is performed by three human experts. These experts use the Leader-1 and Leader-2 robot arms to demonstrate the manipulation tasks, with their actions being synchronously replicated on the Follower-1 and Follower-2 arms. This structured approach ensures that every trajectory is unique, capturing a broad spectrum of challenges and demands encountered in real-world bimanual-mobile robot manipulation.
@@ 3.3 Properties of BRMData
BRMData is structured around diverse manipulation tasks. For detailed task definitions, step-by-step descriptions, please refer to Appendix D.
Table 2: Comparative analysis of robot manipulation tasks across key attributes
@@
3.3.1 Task Diversity
BRMData incorporates a broad spectrum of robotic manipulation tasks that extend from simple tabletop single-arm manipulations to intricate bimanual and mobile manipulations, as detailed in Table 2. The dataset provides a foundation for investigating fundamental robotic interactions like Bottle Pick, Plate Place, and Single Fruit Pick. These tasks encompass basic object handling as well as advanced human-robot interactions. Tasks such as Bottle Handoff and Wine Wipe further the complexity by necessitating coordinated bimanual actions and managing multi-object interactions, representing advanced manipulation challenges. The mobile manipulation tasks, including Garbage Recycle and High Five, introduce navigation and dynamic interaction capabilities, essential for operational efficacy in changing environments. The diverse task range in BRMData enables comprehensive testing and refinement of robotic manipulation strategies suitable for a variety of applications.
@@ 3.3.2 Progressive Difficulty
Single vs. Multi-Object Manipulation: Initially, tasks such as Bottle Pick and Single Fruit Pick involve interactions with a single object, allowing for focused assessments of basic manipulation capabilities. However, as the complexity level rises, the tasks evolve to include multi-object interactions. Tasks like Multi Fruit Pick and Wine Wipe require the robot to manage multiple objects either sequentially or simultaneously. This shift from single to multi-object manipulation tests the robot’s capability in task planning, object prioritization, and execution of complex sequences involving various objects.
Handling Rigid and Flexible Objects: The BRMData dataset differentiates between rigid and flexible objects. Initial tasks like Bottle Pick, Single Fruit Pick, and Plate Place involve rigid objects, which are generally simpler to manipulate due to their stable and predictable nature. These tasks serve as fundamental tests of a robot’s capabilities in gripping, object recognition, and precise placement, laying the groundwork for more complex manipulations. As the tasks progress in complexity, the dataset introduces flexible objects, as seen in Wine Wipe and High Five tasks. The manipulation of flexible objects poses additional challenges due to their deformable nature, which requires more sophisticated sensory and feedback mechanisms to adjust grip strength and manipulation strategies dynamically. For instance, in the Wine Wipe task, the robot must not only lift a wine glass but also clean a spill with a cloth, involving precise movements and pressure control to avoid slippage or damage. The High Five task exemplifies the complexity of interacting with human hands, which introduces unpredictability in object behavior and demands a nuanced understanding of dynamic material properties. This progression from rigid to flexible object manipulation significantly raises the difficulty level, challenging the robot’s capacity to handle a diverse range of material properties and interaction scenarios.
3.3.3 Variety of Collection Strategies
@@
Frame Rate Variability: The dataset incorporates cameras operating at different frame rates to tailor data acquisition to the specific dynamics of each task. For tasks that involve rapid movements, high-frame-rate cameras are utilized to capture fine temporal details, which are crucial for analyzing swift interactions. Conversely, tasksis characterized by slower or more static actions are adequately documented using cameras with standard frame rates. This approach ensures that the temporal resolution of the captured data is precisely aligned with the specific requirements of each task.
3.4 Evaluation Metrics
Robot manipulation tasks are often evaluated using the success rate (SR). While SR provides a measure of task completion success, it does not account for the efficiency of task execution. High SR may still be accompanied by long execution times, which is impractical for real-world applications where time efficiency is crucial. Therefore, we introduce a novel metric: the Manipulation Efficiency Score (MES). It is defined as $\text{MES}=100\frac{\text{SR}}{T}$, where $T$ denotes the average execution time. This metric captures both the effectiveness and efficiency of robot manipulation methods, rewarding those that achieve high success rates with minimal execution time, thereby highlighting both the quality and speed of task execution.

### v2

@@ Abstract
The advancements in embodied AI are increasingly enabling robots to tackle complex real-world tasks, such as household manipulation. However, the deployment of robots in these environments remains constrained by the lack of comprehensive bimanual-mobile robot manipulation data that can be learned. Existing datasets predominantly focus on single-arm manipulation tasks, while the few dual-arm datasets available often lack mobility features, task diversity, comprehensive sensor data, and robust evaluation metrics; they fail to capture the intricate and dynamic nature of household manipulation tasks that bimanual-mobile robots are expected to perform. To overcome these limitations, we propose BRMData, a Bimanual-mobile Robot Manipulation Dataset specifically designed for household applications. BRMData encompasses 10 diverse household tasks, including single-arm and dual-arm tasks, as well as both tabletop and mobile manipulations, utilizing multi-view and depth-sensing data information. Moreover, BRMData features tasks of increasing difficulty, ranging from single-object to multi-object grasping, non-interactive to human-robot interactive scenarios, and rigid-object to flexible-object manipulation, closely simulating real-world household applications. Additionally, we introduce a novel Manipulation Efficiency Score (MES) metric to evaluate both the precision and efficiency of robot manipulation methods in household tasks. We thoroughly evaluate and analyze the performance of advanced robot manipulation learning methods using our BRMData, aiming to drive the development of bimanual-mobile robot manipulation technologies. The dataset is now open-sourced and available at https://embodiedrobot.github.io/.
@@ 1 Introduction
Recent advancements in large language models (LLMs) [1, 2, 3, 4] and multimodal large language models (MLLMs) [5, 6, 7] have significantly advanced the field of embodied manipulation, enhancing the capacity for robots to interact with the 3D physical world in a more human-like manner. This technology aims to endow robots with the ability to perform manipulation tasks, mirroring human dexterity and adaptability. In robot manipulation tasks, household environments recently attract significant attention from researchers due to the prevalence of human-operated tasks and the broad range of practical applications for robots [8, 9]. Moreover, household tasks present a substantial challenge due to their diversity and the complex interactions they entail with various objects. To enable robots to autonomously and efficiently complete household tasks, it is crucial to train manipulation policies for bimanual-mobile robots that combine the manipulation of humanoid robots with autonomous mobility. A central ingredient for training such manipulation policies is diverse training datasets of bimanual-mobile robots.
Recently, related studies have shown that comprehensive robot manipulation datasets are essential for developing and refining policy methods, enabling robots to perform complex household tasks with human-like dexterity [10, 11, 12, 13]. However, existing datasets primarily focus on single-arm manipulation tasks, such as RT-1 [14], RH20T [15], and RT-X [10], which are limited to basic object grasping and placement. This narrow focus fails to address the complexities and coordination demands of bimanual-mobile manipulations for household tasks. Compared with single-arm manipulation datasets, dual-arm manipulation datasets are rare. Although a few datasets include dual-arm manipulations, they often lack mobility features [16] or are limited to simple sensory setups [13], lacking comprehensive sensor data, task variety, progressive task difficulty, and robust evaluation metrics. These restrictions hamper their ability to comprehensively represent the complex tasks that bimanual-mobile robots must perform in diverse household environments.
@@
Based on the aforementioned discussions, we propose a bimanual-mobile Robot Manipulation Dataset (BRMData). BRMData is a novel dataset meticulously designed to advance the development and evaluation of bimanual-mobile robots for household tasks. BRMData captures the intricacies of real-world tasks through human demonstrations, offering a true reflection of the dexterity and interaction complexity required. As depicted in Figure 1, this dataset comprises 10 diverse manipulation tasks, including single-object pick-and-place manipulation, dual-arm coordination, human-robot interactions, and mobile manipulation. These tasks are specifically selected to challenge and develop robot manipulation capabilities incrementally. Moreover, the key properties of BRMData are emphasized as follows:
Task Diversity: BRMData encompasses a broad spectrum of manipulation tasks, including single-arm and dual-arm tasks, as well as tabletop and mobile manipulations. This diversity enhances the adaptability of robot manipulation learning methods across various real-world scenarios.
Progressive Difficulty: BRMData features tasks with escalating complexity, from single-object to multi-object grasping, non-interactive to human-robot interactive scenarios, and rigid-object to flexible-object manipulation. This progression facilitates the development of diverse robot manipulation learning methods.
Sensory Data Comprehensiveness: Employing RGBD camera sensors to perceive multi-view, static and dynamic imaging, depth-sensing and cross-sensor data, BRMData provides comprehensive and essential sensory information for testing the robustness of robot manipulation learning methods.
Manipulation Efficiency Score: We design a novel evaluation metric: Manipulation Efficiency Score (MES), which considers both the success rate and the efficiency of robot manipulation learning methods. This metric provides a more comprehensive evaluation, ensuring that developed manipulation algorithms not only complete tasks effectively but also promptly.
Besides, based on MES, we further thoroughly evaluate the performance of the state-of-the-art robot manipulation learning methods through our BRMData, for promoting the iterative development of the methods.
@@ 2 Related Work
2.1 Robot Manipulation Learning Methods
In the realm of robot learning, the utilization of human demonstrations [17, 18, 19, 20, 14] as a means to imbue robots with manipulation capabilities has garnered significant attention. Robomimic [21] conducts a comprehensive examination of six offline learning algorithms across various simulated and real-world manipulation tasks, delves into the nuances of learning from offline human demonstrations for robot manipulation. DayDreamer [22] enables robots to learn from minimal interaction by planning within a learned world model, which significantly reduces the trial-and-error process typically required in physical settings. CALVIN [23] offers environments for policy learning under language instructions for long-horizon robotic manipulation tasks, emphasizing the importance of comprehensive datasets in enhancing robot-environment interactions. Similarly, DROID [12] provides extensive data for robot manipulation, underscoring the necessity of testing and optimizing algorithms for dynamic, real-world applications. Moreover, ARMBench [24] and PLEX [25] have expanded the application scope by incorporating not only basic manipulation tasks but also more complex scenarios and diverse operational strategies. In addition, action chunking with transformers (ACT) [16] learns a generative model over action sequences based on a conditional variational autoencoder [26]. Diffusion policy [27] is a novel approach to generating robot behavior by representing a robot’s visuomotor policy as a conditional denoising diffusion process.
@@ 2.2 Robot Manipulation Datasets
3 BRMData
In this section, we comprehensively introduce the proposed dataset, called Bimanual-mobile Robot Manipulation Dataset (BRMData), to the community. Firstly, the data collection setup is presented. Then, the properties of BRMData are emphasized. Finally, dataset discussions are given.
3.1 Data Collection Setup
1) Robot Platform: As illustrated in Figure 15, the data collection hardware platform for BRMData comprises 4 ARX-5 robot arms, each equipped with 7 joints and a parallel gripper. The Follower-1 and Follower-2 robot arms, positioned at the front of the platform, are specifically engineered for advanced bimanual manipulations. Each of both arms features a movable RGBD camera sensor at the wrist. The flexibility afforded by these cameras facilitates variable viewing angles and provides detailed visual feedback, crucial for precision in manipulation tasks. Positioned centrally between Follower-1 and Follower-2 is a fixed RGBD camera sensor. This middle camera provides a consistent and broad visual reference across the manipulation area.
Figure 2: Illustration of robot platform.
The Leader-1 and Leader-2 robot arms, located at the rear of the platform, are designated for teleoperation by human experts. This configuration facilitates the synchronous replication of human-directed actions from the leader arms to the follower arms. The platform includes an automated guided vehicle equipped with a two-wheel differential Tracer mobile base for providing essential mobility. The computational demands of the platform are managed by an industrial computer equipped with an NVIDIA 4090 graphics card, an Intel i7-13700 CPU, and 32GB of RAM. This computing setup ensures efficient camera data processing and supports real-time algorithms necessary for effective manipulation and interaction.
2) Data Collection Protocol: The data collection protocol for BRMData is meticulously designed to capture a wide array of manipulation tasks comprehensively. Specifically, as shown in Figure 3, both the left and right wrist-mounted RGBD cameras and the middle camera record RGB and depth images at a resolution of $640\times 480$. The integration of these three views provides a multi-perspective visual input, essential for constructing a detailed 3D representation of the operational environment. Meanwhile, to bring in cross-sensor data, for Bottle Pick and Bottle Handoff tasks, image data are captured using Orbbec Dabai cameras at 30Hz. For other manipulation tasks, Intel Realsense D435 cameras are used to capture image data at 60Hz. In addition, each manipulation task within the dataset includes the collection of 50 action trajectories, gathered at intervals that are specifically tailored to the complexity and specific requirements of each task scenario. To enrich the behavior diversity of the dataset, each manipulation task is performed by three human experts. These experts use the Leader-1 and Leader-2 robot arms to demonstrate the manipulation tasks, with their actions being synchronously replicated on the Follower-1 and Follower-2 arms. This structured approach ensures that every trajectory is unique, capturing a broad spectrum of challenges and demands encountered in real-world bimanual-mobile robot manipulation.
@@ 3.2 Properties of BRMData
BRMData is structured around diverse manipulation tasks. For detailed task definitions and step-by-step descriptions, Please refer to Appendix D. Besides, we further emphasize the following properties of our BRMData.
3.2.1 Task Diversity
BRMData incorporates a broad spectrum of robotic manipulation tasks that extend from single-arm to dual-arm manipulations, and tabletop to mobile manipulations, as detailed in Table 1. The dataset provides a foundation for investigating fundamental robotic interactions like Bottle Pick, Plate Place, and Single Fruit Pick. These tasks encompass basic object handling as well as advanced human-robot interactions. Tasks such as Bottle Handoff and Wine Wipe further the complexity by necessitating coordinated bimanual actions and managing multi-object interactions, representing advanced manipulation challenges. The mobile manipulation tasks, including Garbage Recycle and High Five, introduce navigation and dynamic interaction capabilities, essential for operational efficacy in changing environments. The diverse task range in BRMData enables comprehensive testing and refinement of robotic manipulation strategies suitable for a variety of applications.
Table 1: Key attribute analysis of robot manipulation tasks across BRMData
@@ 3.2.2 Progressive Difficulty
From Single-Object to Multi-Object: Initially, tasks such as Bottle Pick and Single Fruit Pick involve interactions with a single object, allowing for focused assessments of basic manipulation capabilities. However, as the complexity level rises, the tasks evolve to include multi-object interactions. Tasks like Multi Fruit Pick and Wine Wipe require the robot to manage multiple objects either sequentially or simultaneously. This shift from single to multi-object manipulation tests the robot’s capability in task planning, object prioritization, and execution of complex sequences involving various objects.
From Non-Interaction to Human-Robot Interaction: BRMData includes tasks without or with human-robot interactions, as shown in Table 1. Compared with tasks without human-robot interactions, tasks with human-robot interactions (e.g., Fruit Handover and High Five) require a robot to cooperate with humans to achieve the tasks. From the robot perception perspective, human actions are usually different every time, despite the same behavior. This adds a certain amount of random disturbance to the task itself, increasing the difficulty of the task.
From Rigid-Object to Flexible-Object: BRMData differentiates between rigid and flexible objects. Initial tasks like Bottle Pick, Single Fruit Pick, and Plate Place involve rigid objects, which are generally simpler to manipulate due to their stable and predictable nature. These tasks serve as fundamental tests of a robot’s capabilities in gripping, object recognition, and precise placement, laying the groundwork for more complex manipulations. As the tasks progress in complexity, the dataset introduces flexible objects, as seen in Wine Wipe and High Five tasks. The manipulation of flexible objects poses additional challenges due to their deformable nature, which requires more sophisticated sensory and feedback mechanisms to adjust grip strength and manipulation strategies dynamically. For instance, in the Wine Wipe task, the robot must not only lift a wine glass but also clean a spill with a cloth, involving precise movements and pressure control to avoid slippage or damage. The High Five task exemplifies the complexity of interacting with human hands, which introduces unpredictability in object behavior and demands a nuanced understanding of dynamic material properties. This progression from rigid to flexible object manipulation significantly raises the difficulty level, challenging the robot’s capacity to handle a diverse range of material properties and interaction scenarios.
3.2.3 Sensory Data Comprehensiveness
@@
Cross-Sensor: Our BRMData incorporates data collected by different types of cameras. For specific tasks such as the Bottle Pick and Bottle Handoff, image data are captured using Orbbec Dabai cameras operating at a frequency of 30Hz. For more dynamic manipulation tasks, Intel Realsense D435 cameras, which capture data at a higher frequency of 60Hz, are employed to accommodate the quick movements and frequent data updates required. These cross-sensor data enrich our dataset and provide a possibility to test the adaptability of robot manipulation learning methods to different sensing devices.
3.2.4 Manipulation Efficiency Score
Evaluation metrics are crucial to measuring the performance of the robot manipulation learning methods. Success rate (SR) is generally used to evaluate the performance of the methods in robot manipulation tasks. While SR provides a measure of task completion success, it does not account for the efficiency of task execution. High SR may still be accompanied by long execution times, which is impractical for real-world applications where time efficiency is crucial. Therefore, we introduce a novel metric: Manipulation Efficiency Score (MES). It is defined as $\text{MES}=100\frac{\text{SR}}{T}$, where $T$ denotes the average execution time. This metric captures both the effectiveness and efficiency of robot manipulation methods, rewarding those that achieve high success rates with minimal execution time, thereby highlighting both the quality and speed of task execution.
3.3 Dataset Discussions
To further analyze the highlights of our BRMData, we compare and discuss our dataset with existing datasets, as shown in Table 2. According to the comparison, BRMData stands out due to its comprehensive coverage of bimanual-mobile manipulations. Through the integration of dual-arm coordination, diverse task settings, and comprehensive sensor data collection, BRMData offers an enriched training environment unparalleled by existing representative datasets. Specifically, compared with the other datasets, BRMData incorporates not only human-robot tasks but also cross-sensor tasks, along with depth-sensing information. This enhancement mitigates previous limitations and empowers robots to operate with heightened versatility and adaptability, particularly in dynamically changing settings. In addition, BRMData has more unique properties than the other datasets: task diversity, progressive difficulty, and sensory data comprehensiveness, which can greatly facilitate research in robot manipulation learning.
Table 2: Comparison to existing datasets for robot manipulation.
Dataset Human Demonstration Movable Dual-arm Depth-sensing Multi-view Torque Human-Robot Interaction Cross-Sensor
MIME [17] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
RT-1 [14] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
RH20T [15] $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
BridgeData V2 [29] $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
ALOHA [16] $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$
Mobile ALOHA [13] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
BRMData (Ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$